OpenStudy (chrisplusian):

Suppose that u and v are orthogonal vectors in R^n. Find the value of the expression <3u-v. 3u+v>.

OpenStudy (chrisplusian):

The notation used here "<x,y>" is the way they represent the dot product of vector x and vector y in my current textbook

OpenStudy (irishboy123):

you can distribute in the dot product so use FOIL and then bear in mind that \(\mathbf u \bullet \mathbf v = 0 \) and vice versa

OpenStudy (irishboy123):

if that worries you, start with this idea and expand it out to 4 vectors \[ \mathbf A \cdot ( \mathbf B + \mathbf C ) = \mathbf A \cdot \mathbf B + \mathbf A \cdot \mathbf C .\] now let \(\mathbf A = \mathbf D + \mathbf E \)